rvc (rvc):

the ph of 10^-8M HCl is A less than 7 B 8 C greater than 7 D 7

OpenStudy (ikram002p):

less , HCL ph is always less than 7 cuz its acid

rvc (rvc):

one more the concentration of OH^- is 1 x 10^-10 the solution is A acidic B neutral C basic D none of the above

rvc (rvc):

@ikram002p @thomaster

OpenStudy (ikram002p):

will , is the OH in water solution ? if yes then the concentration dnt matter OH is always basic so i cant juge according to the concentration only i need to know what is the solute

rvc (rvc):

solute is not given in the question that's y i have a doubt

rvc (rvc):

@ikram002p

rvc (rvc):

@Jack1

OpenStudy (ikram002p):

humm then sry i cant help :O but i think im lost btw C ( usually water ) and D ( cuz its not given extra info )

rvc (rvc):

the answer given is acidic @Preetha @theEric

rvc (rvc):

can u help me jack?

rvc (rvc):

hey?

OpenStudy (jack1):

been a while since i did chem, but think this is correct: acid is a hydrogen donor base is a hydrogen acceptor so OH is ...hydroxide? and should therefore be a base... but this is OH^-... so dont recall and what is the standard ph of water? (i think it's pretty close to 7, so maybe it doesnt affect the outcome also: a concentration of 1x 10^(-10) is that strong or weak?

OpenStudy (jack1):

so overall i would have thought "basic"

rvc (rvc):

according to lewis theory acid:which accets a share of e base : donates a share of e

rvc (rvc):

@Jack1

rvc (rvc):

@Preetha @sarah786 @random231

OpenStudy (anonymous):

You have been given some incorrect information. \[\text{K}_w = [\text{H}^{1+}][\text{OH}^{1-}]\] \[\text{If} ~ \text{K}_w = 1.0 \times 10^{-14}\] \[[\text{OH}^{1-}] = \frac{\text{K}_w}{[\text{H}^{1+}]}\text{and} ~[\text{H}^{1+}] = \frac{\text{K}_w}{[\text{OH}^{1-}]}\\ \text{pH} = - \log [\text{H}^{1+}]\] See if you can take it from here.

OpenStudy (vincent-lyon.fr):

@ghuczek is right. You must follow his steps. (Some of us misread your 2nd question, because it looked like the first one)